### v2

# Lattice point enumeration of polytopes associated to integer compositions

Abstract.
An $n$-dimensional lattice polytope ${\mathcal{Q}}_{\sigma}$ can be associated to any composition $\sigma$ of a positive integer $n$, as a special case of constructions due to Pitman–Stanley and Chapoton. The entries of the $h$-vector of $\sigma$, introduced by Chapoton, enumerate the lattice points in ${\mathcal{Q}}_{\sigma}$ by the number of their zero coordinates. Chapoton conjectured that this vector is equal to the $h$-vector of a simplicial polytope. This paper proves that the $h$-vector of $\sigma$ is gamma-positive, hence palindromic and unimodal, and confirms certain other conjectures of Chapoton on the lattice point enumeration of composition polytopes. A combinatorial interpretation of their $h^{\ast}$-polynomials is deduced.
Mathematics Subject Classifications: 52B20, 05A15, 05E45, 06A07
Key words and phrases. Lattice polytope, integer composition, Ehrhart polynomial, poset, zeta polynomial, gamma-positivity.
1. Introduction
A composition of a positive integer $n$ is defined as a sequence $\sigma=(r_{1},r_{2},\dots,r_{k})$ of positive integers whose sum is equal to $n$. Given such a composition $\sigma$, we consider the polytope ${\mathcal{Q}}_{\sigma}$ in ${\mathbb{R}}^{n}$ defined by the inequalities $x_{i}\geq 0$ for $1\leq i\leq n$ and
(1) $x_{1}+x_{2}+\cdots+x_{s_{i}}\leq s_{i}$
for $1\leq i\leq k$, where $s_{i}=r_{1}+r_{2}+\cdots+r_{i}$ and $x_{1},x_{2},\dots,x_{n}$ are the standard coordinates of ${\mathbb{R}}^{n}$. We refer to the polytopes ${\mathcal{Q}}_{\sigma}$ in this paper as composition polytopes.
Composition polytopes are special cases of the polytopes associated to arbors (defined as rooted trees on the vertex set of blocks of a partition of a finite set), introduced and studied by Chapoton [5]. They also arise as Pitman–Stanley polytopes [12] when some of the parameters defining these polytopes are allowed to be zero. The composition polytope ${\mathcal{Q}}_{\sigma}$ reduces to the $n$th dilate of the $n$-dimensional simplex in ${\mathbb{R}}^{n}$ with vertices the origin and the unit coordinate vectors, when $\sigma$ has one part, and to the Pitman–Stanley polytope with all parameters equal to 1, when all parts of $\sigma$ are equal to 1.
The face lattice of composition polytopes is well understood. As observed in [12, Theorem 20], ${\mathcal{Q}}_{\sigma}$ is combinatorially isomorphic to a product of simplices, the dimensions of which are the parts of $\sigma$. The lattice point enumeration of composition polytopes, which is more interesting, has been studied for the more general classes of arbor polytopes [5] and Pitman–Stanley polytopes [12]. Chapoton defined the $h$-vector of a composition $\sigma$ of $n$ (more generally, of an arbor) as the sequence $h(\sigma)=(h_{i}(\sigma))_{0\leq i\leq n}$, where $h_{i}(\sigma)$ is the number of points in ${\mathcal{Q}}_{\sigma}\cap{\mathbb{N}}^{n}$ having exactly $i$ nonzero coordinates (this vector should not be confused with the $h$-vector of the simple polytope ${\mathcal{Q}}_{\sigma}$). We note that $h_{0}(\sigma)=h_{n}(\sigma)=1$. The following intriguing conjecture has been stated more generally for the class of arbor polytopes [5, Conjecture 0.1]; we refer to the discussion in [5, Section 0.1] for the motivation behind it.
Conjecture 1.1 ([5]).
The vector $h(\sigma)$ is equal to the $h$-vector of an $n$-dimensional simplicial polytope for every composition $\sigma$ of $n$.
In particular, $h(\sigma)$ should be palindromic and unimodal, meaning that $h_{i}(\sigma)=h_{n-i}(\sigma)$ for all $0\leq i\leq n$ and $h_{0}(\sigma)\leq h_{1}(\sigma)\leq\cdots\leq h_{\lfloor n/2\rfloor}(\sigma)$. Moreover, it is suggested in [5, Section 0.1] that $h(\sigma)$ should in fact be equal to the $h$-vector of an $n$-dimensional flag simplicial polytope. Thus, in view of Gal’s conjecture [6, Conjecture 2.1.7] (see also [1, Section 3]) on gamma-positivity, it is reasonable to expect that the generating polynomial $h(\sigma,t):=\sum_{i=0}^{n}h_{i}(\sigma)t^{i}$ of $h(\sigma)$ is $\gamma$-positive, meaning that
(2) $h(\sigma,t)=\sum_{i=0}^{\lfloor n/2\rfloor}\gamma_{i}(\sigma)t^{i}(1+t)^{n-2i}$
for some nonnegative integers $\gamma_{i}(\sigma)$. The first main result of this paper shows that this is indeed the case.
Theorem 1.2.
The polynomial $h(\sigma,t)$ is $\gamma$-positive for every composition $\sigma$. In particular, the vector $h(\sigma)$ is palindromic and unimodal.
Our second main result confirms another curious conjecture of Chapoton [5], concerning the Ehrhart polynomial of composition polytopes. We recall (see, for instance, [10, Section 4.6]) that the Ehrhart polynomial ${\rm Ehr}(Q,t)$ of a lattice polytope ${\mathcal{Q}}\subset{\mathbb{R}}_{\geq 0}^{n}$ has the defining property that ${\rm Ehr}({\mathcal{Q}},m)=\#(m{\mathcal{Q}}\cap{\mathbb{N}}^{n})$ for every $m\in{\mathbb{N}}$. The zeta polynomial (see [10, Section 3.12]) ${\mathcal{Z}}(P,t)$ of a finite poset $(P,\preceq)$ has the defining property that ${\mathcal{Z}}(P,m+1)$ is equal to the number of $m$-element multichains in $P$, meaning sequences $(p_{1},p_{2},\dots,p_{m})$ of elements of $P$ such that $p_{1}\preceq p_{2}\preceq\cdots\preceq p_{m}$, for every $m\in{\mathbb{N}}$. Given a composition $\sigma$ of $n$, Chapoton [5, Section 1] has considered the set $P_{\sigma}$ of lattice points in ${\mathcal{Q}}_{\sigma}$ as a subposet of the product poset ${\mathbb{N}}^{n}$.
Theorem 1.3.
([5, Conjecture 3.1]) We have ${\rm Ehr}({\mathcal{Q}}_{\sigma},t)={\mathcal{Z}}(P_{{\rm rev}(\sigma)},t+1)$ for every composition $\sigma$, where ${\rm rev}(\sigma)$ is the composition obtained from $\sigma$ by reversing the order in which the parts of $\sigma$ appear.
Here is another intriguing conjecture from [5], stated there for all arbor polytopes (see [5, Conjecture 2.2]), on which no progress is made in this paper.
Conjecture 1.4 ([5]).
For every composition $\sigma$, all roots of the Ehrhart polynomial of ${\mathcal{Q}}_{\sigma}$ are real and lie in the interval [1,0][-1,0].
Theorem 1.2 is proven in Section 2; the proof yields an explicit combinatorial interpretation of the numbers $\gamma_{i}(\sigma)$ which appear in Equation (2) (see Proposition 2.2). Section 2 also shows that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma$. Two proofs of Theorem 1.3, one essentially bijective and another based on the results of [12, Section 3], are given in Section 3 and a combinatorial interpretation of the $h^{\ast}$-polynomial of ${\mathcal{Q}}_{\sigma}$ is deduced via the theory of pure shellability of posets (see Proposition 3.2). While this paper is restricted to composition polytopes, we do hope that some of the arguments presented here can be extended to more general classes of lattice polytopes.
Throughout this paper, we set ${\mathbb{N}}=\{0,1,2,\dots\}$ and $n]=\{1,2,\dots,n\}$ for $n\in{\mathbb{N}}$. We denote by #S\#S the cardinality of a finite set $S$ and write $\sigma\models n$ when $\sigma$ is a composition of $n$. Our notation and terminology about posets follows [10, Chapter 3]. Background on Ehrhart theory and face enumeration of simplicial complexes and polytopes can be found in [3, 7] [10, Chapter 4] and [9], respectively.
2. Gamma-positivity
This section proves Theorem 1.2 and shows (see Proposition 2.5) that $h(\sigma)=h({\rm rev}(\sigma))$ for every composition $\sigma$.
Throughout this section we let $\sigma=(r_{1},r_{2},\dots,r_{k})$ be a composition of $n$ and set $s_{i}=r_{1}+r_{2}+\cdots+r_{i}$ for $i\in[k]$. We recall that $P_{\sigma}$ consists of all points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that
$a_{1}+a_{2}+\cdots+a_{s_{i}}\leq s_{i}$
for every $i\in[k]$. We record such a point by the $n$-tuple $(p_{1},p_{2},\dots,p_{n})\in{\mathbb{N}}^{n}$, where
[…]
for every $i\in[n]$. As a running example, if $\sigma=(3,4,2)$, then $(0,0,2,0,4,0,0,3,0)\in P_{\sigma}$ is recorded by the tuple $(0,0,2,0,6,0,0,9,0)$. The map just described shows that $P_{\sigma}$ is in one-to-one correspondence with the set, say ${\mathcal{H}}_{\sigma}$, of $n$-tuples $(p_{1},p_{2},\dots,p_{n})\in{\mathbb{N}}^{n}$ consisting of positive integers which appear in strictly increasing order from left to right and zeros, such that $p_{1},p_{2},\dots,p_{s_{i}}\leq s_{i}$ for every $i\in[k]$. Moreover, $h_{j}(\sigma)$ is equal to the number of elements of ${\mathcal{H}}_{\sigma}$ which have exactly $j$ nonzero entries.
Given $x\in{\mathcal{H}}_{\sigma}$, we denote by $S(x)$ the set of positions of the nonzero entries in $x$ and by $T(x)$ the set of these entries. Thus, in our running example, we have $S(x)=\{3,5,8\}$ and $T(x)=\{2,6,9\}$. Clearly, $S(x),T(x)\subseteq[n]$ and #S(x)=#T(x)\#S(x)=\#T(x) for every $x\in{\mathcal{H}}_{\sigma}$. Moreover, every $x\in{\mathcal{H}}_{\sigma}$ is determined by $S(x)$ and $T(x)$ and
(3) $h(\sigma,t)=\sum_{x\in{\mathcal{H}}_{\sigma}}t^{\#S(x)}.$
Lemma 2.1.
Let $S,T$ be subsets of $[n]$ of equal cardinalities and let $R\subseteq[n]$ be disjoint from $S\cup T$. Then, the pair $(S,T)$ has the property that $S=S(x)$ and $T=T(x)$ for some $x\in{\mathcal{H}}_{\sigma}$ if and only if so does $(S\cup R,T\cup R)$.
Proof.
We may assume that $R$ has a single element $r\in[n]$. Let $x=(p_{1},p_{2},\dots,p_{n})\in{\mathbb{N}}^{n}$ be the $n$-tuple which has the elements of $T$ as its nonzero entries and these appear in strictly increasing order from left to right in the positions of the elements of $S$. Similarly, let $y=(q_{1},q_{2},\dots,q_{n})\in{\mathbb{N}}^{n}$ be the $n$-tuple which has the elements of $T\cup\{r\}$ as its nonzero entries and these appear in strictly increasing order from left to right in the positions of the elements of $S\cup\{r\}$. We have $x\in{\mathcal{H}}_{\sigma}$ if and only if
(4) $p_{1},p_{2},\dots,p_{s_{i}}\leq s_{i}\ \ \text{for every}\ i\in[k]$
and $y\in{\mathcal{H}}_{\sigma}$ if and only if
(5) $q_{1},q_{2},\dots,q_{s_{i}}\leq s_{i}\ \ \text{for every}\ i\in[k].$
Thus, we have to show the equivalence of (4) and (5). This is obvious if $q_{r}=r$, since then $x$ and $y$ differ only in the $r$th coordinate. Suppose that $r=q_{s}$ for some $s>r$. Then, $y$ is obtained from $x$ by placing $r$ in position $s$ and shifting to the left some nonzero entries, all less than $r$, within positions $r,r+1,\dots,s$. Finally, if $r=q_{s}$ for some $s<r$, then $y$ is obtained from $x$ by placing $r$ in position $s$ and shifting to the right some nonzero entries, all larger than $r$, within positions $s,s+1,\dots,r$. The equivalence (4) $\Leftrightarrow$ (5) holds in either case. ∎
Theorem 1.2 is a consequence of the following result.
Proposition 2.2.
Equation (2) holds if $\gamma_{i}(\sigma)$ is defined as the number of elements $x\in{\mathcal{H}}_{\sigma}$ such that $S(x)\cap T(x)=\varnothing$ and #S(x)=i\#S(x)=i, for $0\leq i\leq\lfloor n/2\rfloor$.
Proof.
For $x,y\in{\mathcal{H}}_{\sigma}$ let us write $x\sim_{\sigma}y$ if
$S(x){\smallsetminus}(S(x)\cap T(x))$ $=S(y){\smallsetminus}(S(y)\cap T(y))$
$T(x){\smallsetminus}(S(x)\cap T(x))$ $=T(y){\smallsetminus}(S(y)\cap T(y)).$
Clearly, $\sim_{\sigma}$ is an equivalence relation on ${\mathcal{H}}_{\sigma}$. In view of Lemma 2.1, every equivalence class has a unique representative $x\in{\mathcal{H}}_{\sigma}$ such that $S(x)\cap T(x)=\varnothing$ and the map which assigns the pair $(S(y),T(y))$ to $y\in{\mathcal{H}}_{\sigma}$ induces a bijection from the equivalence class, say $C_{x}$, of $x$ to the set of pairs $(S(x)\cup R,T(x)\cup R)$, where $R$ ranges over all subsets of $n]{\smallsetminus}(S(x)\cup T(x))$. As a result,
$\sum_{y\in C_{x}}t^{\#S(y)}=\sum_{R\subseteq[n]{\smallsetminus}(S(x)\cup T(x))}t^{\#(S(x)\cup R)}=t^{\#S(x)}(1+t)^{n-2\,\#S(x)}.$
We sum over all equivalence classes of $\sim_{\sigma}$ and the proof follows from (3). ∎
Example 2.3.
Let $\sigma=(n)$ be the composition of $n$ with only one part. Then, $P_{\sigma}$ consists of all points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that $a_{1}+a_{2}+\cdots+a_{n}\leq n$, $h_{i}(\sigma)$ is equal to the total number of pairs $(S,T)$ of $i$-element subsets of $[n]$ and $\gamma_{i}(\sigma)$ is equal to the number of such pairs $(S,T)$ with $S\cap T=\varnothing$. Hence, $h_{i}(\sigma)={n\choose i}^{2}$ for $0\leq i\leq n$ and $\gamma_{i}(\sigma)={n\choose i}{n-i\choose i}$ for $0\leq i\leq\lfloor n/2\rfloor$ and Proposition 2.2 specializes to the well known identity
$\sum_{i=0}^{n}{n\choose i}^{2}t^{i}=\sum_{i=0}^{\lfloor n/2\rfloor}{n\choose i}{n-i\choose i}t^{i}(1+t)^{n-2i}$
(see [1, Theorem 2.32] for a generalization in the setting of finite Coxeter groups). ∎
For the remainder of this section we represent elements of ${\mathbb{N}}^{n}$ with sum of coordinates at most $n$ by lattice paths in the $xy$-plane as follows. A point $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ with $a_{1}+a_{2}+\cdots+a_{n}\leq n$ is represented by the lattice path with initial vertex (1,0)(-1,0), terminal vertex $(n,n+1)$ and east (E) or north (N) unit steps, so that the initial and final steps are east and north, respectively, and the height of the part of the lattice path within the strip $j-1\leq x\leq j$ is $a_{1}+a_{2}+\cdots+a_{j}$ for every $j\in[n]$. We observe that the number of nonzero entries of such an element of ${\mathbb{N}}^{n}$ is equal to one less than the number of east-north corners of the associated lattice path.
We denote by ${\mathcal{G}}_{n}$ the set of lattice paths which represent the elements of ${\mathbb{N}}^{n}$ with sum of coordinates at most $n$ and let $\Gamma_{\sigma}\in{\mathcal{G}}_{n}$ be the lattice path associated to the $n$-tuple with entry $s_{i}$ in position $s_{i-1}+1$ for $1\leq i\leq k$ (where $s_{0}:=0$) and all other entries equal to zero. Thus, $\Gamma_{\sigma}$ has successive steps east, followed by $r_{1}$ north and $r_{1}$ east steps, followed by $r_{2}$ north and $r_{2}$ east steps, and so on, followed by a string of north steps at the end. Then, the elements of $P_{\sigma}$ are represented by the lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below $\Gamma_{\sigma}$.
Example 2.4.
Let $\sigma=(1,1,\dots,1)\models n$, so that $P_{\sigma}$ consists of all points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that $a_{1}+a_{2}+\cdots+a_{j}\leq j$ for every $j\in[n]$. The lattice path $\Gamma_{\sigma}$ has successive steps $\textrm{ENEN}\cdots\textrm{EN}$ and the total number of lattice paths in ${\mathcal{G}}_{n}$ which lie weakly below it (and hence the cardinality of $P_{\sigma}$) is well known to equal the Catalan number $\frac{1}{n+2}{2n+2\choose n+1}$. Moreover, counting these lattice paths by the number of east-north corners shows that $h_{i}(\sigma)$ is equal to the Narayana number
$h_{i}(\sigma)=\frac{1}{i+1}{n\choose i}{n+1\choose i}$
for $0\leq i\leq n$ (see, for instance, [11, Problem A.46]).
By Proposition 2.2, $\gamma_{i}(\sigma)$ is equal to the number of pairs $(S,T)$ of disjoint $i$-element subsets of $[n]$ such that the $r$th largest element of $T$ is smaller that the $r$th largest element of $S$ for every $r$. Since these pairs are in one-to-one correspondence with Young tableaux of shape $(i,i)$ with entries in $[n]$, we have
$\gamma_{i}(\sigma)$ $={n\choose 2i}\cdot\frac{1}{i+1}{2i\choose i}=\frac{1}{i+1}\cdot\frac{n!}{(i!)^{2}\,(n-2i)!}$
$=\frac{1}{i+1}{n\choose i}{n-i\choose i}$
for $0\leq i\leq\lfloor n/2\rfloor$ and Proposition 2.2 specializes to the well known identity
$\sum_{i=0}^{n}\frac{1}{i+1}{n\choose i}{n+1\choose i}t^{i}=\sum_{i=0}^{\lfloor n/2\rfloor}\frac{1}{i+1}{n\choose i}{n-i\choose i}t^{i}(1+t)^{n-2i}$
(see, for instance, again [1, Theorem 2.32]). ∎
We conclude this section with the following statement.
Proposition 2.5.
([5, Conjecture 3.2]) We have $h(\sigma)=h({\rm rev}(\sigma))$ for every $\sigma\models n$.
Proof.
We observe that orthogonal reflection in the line $x+y=n$ maps $\Gamma_{\sigma}$ to $\Gamma_{{\rm rev}(\sigma)}$ and hence induces a bijection from the set of lattice paths which represent the elements of $P_{\sigma}$ to the set of those for $P_{{\rm rev}(\sigma)}$. We also observe that the number of east-north corners of a lattice path is preserved under this bijection and the proof follows. ∎
3. Ehrhart and Zeta polynomials
This section offers two proofs of Theorem 1.2 and deduces a combinatorial interpretation of the $h^{\ast}$-polynomial of the composition polytope ${\mathcal{Q}}_{\sigma}$. As in the previous section, $\sigma=(r_{1},r_{2},\dots,r_{k})$ is a composition of $n$ and $s_{i}=r_{1}+r_{2}+\cdots+r_{i}$ for $0\leq i\leq k$, where $s_{0}:=0$.
We recall that ${\mathbb{N}}^{n}$ is partially ordered as a direct product, meaning that for $x,y\in{\mathbb{N}}^{n}$ with $x=(a_{1},a_{2},\dots,a_{n})$ and $y=(b_{1},b_{2},\dots,b_{n})$ we have $x\preceq y$ if and only if $a_{i}\leq b_{i}$ for every $i\in[n]$, where $\preceq$ stands for the partial order of ${\mathbb{N}}^{n}$.
Lemma 3.1.
There is a one-to-one correspondence from the set of $m$-multichains in ${\mathbb{N}}^{n}$ with a given maximal element $(a_{1},a_{2},\dots,a_{n})$ to the set of solutions $(b_{i,j})_{1\leq i\leq n,\,1\leq j\leq m}\in{\mathbb{N}}^{mn}$ of the system of linear equations
$x_{i,1}+x_{i,2}+\cdots+x_{i,m}=a_{i}$
for $i\in[n]$.
Proof.
By definition of the product order ${\mathbb{N}}^{n}$, it suffices to verify the statement for $n=1$, which is a trivial task. ∎
First proof of Theorem 1.3.
For any $m\in{\mathbb{N}}$, ${\rm Ehr}(Q_{\sigma},m)$ is equal to the number of points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that
(6) $a_{1}+a_{2}+\cdots+a_{s_{i}}\leq s_{i}m$
for $i\in[k]$. Thus, for $i=k$ we get $a_{1}+a_{2}+\cdots+a_{n}\leq mn$. Let’s consider the partial sums $a_{1}+a_{2}+\cdots+a_{i}$ and, for $0\leq j\leq mn$, let’s denote by $b_{j}$ the number of indices $i\in[n]$ for which $a_{1}+a_{2}+\cdots+a_{i}=j$ (the reader may want to interpret the $b_{j}$ when $(a_{1},a_{2},\dots,a_{n})$ is represented by a lattice path, as in Section 2). Then,
(7) $b_{0}+b_{1}+\cdots+b_{mn}=n$
and
(8) $b_{0}+b_{1}+\cdots+b_{s_{i}m}\geq s_{i}$
for $i\in[k]$. Moreover, the map which sends $(a_{1},a_{2},\dots,a_{n})$ to $(b_{0},b_{1},\dots,b_{mn})$ is a one-to-one correspondence from the set of solutions $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ of the system of linear inequalities (6) to the set of solutions $(b_{0},b_{1},\dots,b_{mn})\in{\mathbb{N}}^{mn+1}$ of the system of linear equalities and inequalities (7) and (8). Eliminating $b_{0}$ we find that the latter set is in one-to-one correspondence with the set of solutions $(b_{1},b_{2},\dots,b_{mn})\in{\mathbb{N}}^{mn}$ of the system of linear inequalities
(9) $b_{s_{i}m+1}+b_{s_{i}m+2}+\cdots+b_{mn}\leq n-s_{i}$
for $i\in[k-1]$ and
(10) $b_{1}+b_{2}+\cdots+b_{mn}\leq n.$
The number of these solutions is equal to the number ${\mathcal{Z}}(P_{{\rm rev}(\sigma)},m+1)$ of $m$-multichains in $P_{{\rm rev}(\sigma)}$. This follows from the fact that, since $n-s_{i}=r_{i+1}+r_{i+2}+\cdots+r_{k}$, the set $P_{{\rm rev}(\sigma)}$ consists of all tuples $(x_{1},x_{2},\dots,x_{n})\in{\mathbb{N}}^{n}$ such that
$x_{s_{i}+1}+x_{s_{i}+2}+\cdots+x_{n}\leq n-s_{i}$
for $0\leq i<k$ and from Lemma 3.1. ∎
Part of our second proof of Theorem 1.3 follows the proof of [12, Equation (33)], which is a formula for the number of lattice points of a Pitman–Stanley polytope. To avoid making too long of a digression, we assume familiarity with Sections 3 and 4 of [12] and only sketch this part of the proof.
Second proof of Theorem 1.3.
Let us consider the poset $\Omega_{\sigma}$ which consists of two disjoint chains $\alpha_{1}\prec\alpha_{2}\prec\cdots\prec\alpha_{n}$ and $\beta_{1}\prec\beta_{2}\prec\cdots\prec\beta_{n}$, together with the additional cover relations $\alpha_{s_{i}}\prec\beta_{s_{i}}$ for every $i\in[k]$. We let $C=\{\beta_{1},\beta_{2},\dots,\beta_{n}\}$ and ${\mathbf{u}}=(u_{1},u_{2},\dots,u_{n})$, where $u_{j}=jm$ for every $j\in[n]$ and where $m\in{\mathbb{N}}$ is fixed. In this situation, Pitman and Stanley [12, Section 3] have defined the order cone section ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as the polytope of all order preserving maps $f:\Omega_{\sigma}\rightarrow{\mathbb{R}}_{\geq 0}$ such that $f(\beta_{j})=u_{j}$ for every $j\in[n]$. Setting $f_{j}=f(\alpha_{j})$ for every $j\in[n]$ and $f_{0}=0$ we have
$0=f_{0}\leq f_{1}\leq\cdots\leq f_{n}$
and
$f_{s_{i}}\leq s_{i}m$
for every $i\in[k]$. The transformation $x_{j}=f_{j}-f_{j-1}$ for $j\in[n]$ then shows that ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ is integrally equivalent to $m{\mathcal{Q}}_{\sigma}$. Theorem 5 in [12] expresses the number of lattice points of ${\mathcal{C}}_{C}(\Omega_{\sigma},{\mathbf{u}})$ as a sum over all linear extensions of $\Omega_{\sigma}$. Such a linear extension is uniquely determined by the positions, say $1\leq j_{1}<j_{2}<\cdots<j_{n}\leq 2n$, of $\beta_{1},\beta_{2},\dots,\beta_{n}$, which must satisfy $j_{s_{i}}\geq 2s_{i}$ for every $i\in[k]$. Thus, setting $\kappa_{i}=j_{i}-j_{i-1}-1$ for $i\in[n]$ (where $j_{0}=0$), the argument of [12, p. 617] shows that in our situation, [12, Theorem 5] yields the formula
(11) ${\rm Ehr}({\mathcal{Q}}_{\sigma},m)=\sum_{(\kappa_{1},\kappa_{2},\dots,\kappa_{n})\in K_{\sigma}}{m+\kappa_{1}\choose\kappa_{1}}{m+\kappa_{2}-1\choose\kappa_{2}}\cdots{m+\kappa_{n}-1\choose\kappa_{n}}$,
where $K_{n}$ consists of all tuples $(\kappa_{1},\kappa_{2},\dots,\kappa_{n})\in{\mathbb{N}}^{n}$ such that $\kappa_{1}+\kappa_{2}+\cdots+\kappa_{n}=n$ and $\kappa_{1}+\kappa_{2}+\cdots+\kappa_{s_{i}}\geq s_{i}$ for every $i\in[k]$.
Thus, it suffices to prove that the right-hand side of Equation (11) is also equal to the number ${\mathcal{Z}}(P_{{\rm rev}(\sigma)},m+1)$ of $m$-element multichains in $P_{{\rm rev}(\sigma)}$. As already observed, $P_{{\rm rev}(\sigma)}$ consists of all tuples $(\lambda_{1},\lambda_{2},\dots,\lambda_{n})\in{\mathbb{N}}^{n}$ such that $\lambda_{1}+\lambda_{2}+\cdots+\lambda_{n}\leq n$ and
$\lambda_{s_{i}+1}+\lambda_{s_{i}+2}+\cdots+\lambda_{n}\leq n-s_{i}$
for $i\in[k-1]$. Setting $\lambda_{0}=n-\lambda_{1}-\lambda_{2}-\cdots-\lambda_{n}\geq 0$, we have $\lambda_{0}+\lambda_{1}+\cdots+\lambda_{n}=n$ and
$\lambda_{0}+\lambda_{1}+\cdots+\lambda_{s_{i}}\geq s_{i}$
for every $i\in[k]$ and ${\mathcal{Z}}(P_{{\rm rev}(\sigma)},m+1)$ is equal to the sum of the number of $m$-multichains in ${\mathbb{N}}^{n}$ with maximum element $(\lambda_{1},\lambda_{2},\dots,\lambda_{n})$ over all such tuples $(\lambda_{0},\lambda_{1},\dots,\lambda_{n})\in{\mathbb{N}}^{n+1}$. We may set $\kappa_{1}=\lambda_{0}+\lambda_{1}$ and $\kappa_{i}=\lambda_{i}$ for all $1<i\leq n$ and sum instead the number of $m$-multichains in ${\mathbb{N}}^{n}$ with maximum element $(\lambda_{1},\kappa_{2},\dots,\kappa_{n})$, where $0\leq\lambda_{1}\leq\kappa_{1}$, over all tuples $(\kappa_{1},\kappa_{2},\dots,\kappa_{n})\in K_{n}$. By Lemma 3.1 and basic enumerative combinatorics, the number of these multichains is equal to the right-hand side of Equation (11) and the proof follows. ∎
We now present the final main result of this paper. We recall that the $h^{\ast}$-polynomial of an $n$-dimensional polytope in ${\mathbb{R}}^{n}$ with vertices in ${\mathbb{Z}}^{n}$ is defined by the equation
(12) $\sum_{m\geq 0}{\rm Ehr}({\mathcal{Q}},m)t^{m}=\frac{h^{\ast}({\mathcal{Q}},t)}{(1-x)^{n+1}}.$
Given a word $w=(w_{1},w_{2},\dots,w_{n})\in[n]^{n}$, we denote by ${\rm asc}(w)$ the number of indices (called ascents) $i\in[n]$ such that $w_{i-1}<w_{i}$, where $w_{0}:=1$.
Proposition 3.2.
We have
(13) $h^{\ast}({\mathcal{Q}}_{\sigma},t)=\sum_{w}t^{{\rm asc}(w)},$
where $w$ ranges over all words $w\in[n]^{n}$ which have at least $s_{i}$ entries less than or equal to $s_{i}$ for every $i\in[k]$. Equivalently, the unique nondecreasing rearrangement $(u_{1},u_{2},\dots,u_{n})$ of $w$ should satisfy $u_{s_{i}}\leq s_{i}$ for every $i\in[k]$.
The proof employs the basics of the theory of EL-shellability of posets [4] [13, Lecture 3], which we briefly recall. Let $(P,\preceq)$ be a finite graded poset which has a minimum element $\hat{0}$ and a maximum element $\hat{1}$ and let ${\mathcal{E}}(P)$ be the set of covering relations of $P$. An edge labeling of $P$ (with integer labels) is any map $\lambda:{\mathcal{E}}(P)\rightarrow{\mathbb{Z}}$. Let $C$ be an unrefinable chain $p_{0}\prec p_{1}\prec\cdots\prec p_{m}$ of elements of $P$, so that $(p_{i-1},p_{i})\in{\mathcal{E}}(P)$ for every $i\in[m]$. We define the label of $C$ with respect to $\lambda$ as the sequence
$\lambda(C)=(\lambda(p_{0},p_{1}),\lambda(p_{1},p_{2}),\dots,\lambda(p_{m-1},p_{m}))$
and the descent set of $C$ as ${\rm Des}(C)=\{i\in[m-1]:\lambda(p_{i-1},p_{i})>\lambda(p_{i},p_{i+1})\}$. We say that $C$ is rising with respect to $\lambda$ if $\lambda(C)$ is weakly increasing and that $C$ is lexicographically smaller (with respect to $\lambda$) than an unrefinable chain $C^{\prime}$ in $P$ of the same length if $\lambda(C)$ precedes $\lambda(C^{\prime})$ in the lexicographic order. An edge labeling $\lambda$ of $P$ is called an EL-labeling if for every nonsingleton interval $[u,v]$ in $P$: (i) there is a unique rising (with respect to $\lambda$) maximal chain in $[u,v]$; and (ii) this chain is lexicographically smallest among all maximal chains in $[u,v]$. Then
(14) $h(\Delta(P),t)=\sum_{C\in{\mathcal{M}}(P)}t^{{\rm des}(C)},$
where ${\mathcal{M}}(P)$ is the set of all maximal chains of $P$ and ${\rm des}(C)=\#{\rm Des}(C)$ is the number of descents of $C$ for $C\in{\mathcal{M}}(P)$. Here $\Delta(P)$ stands for the order complex of $P$ (meaning, the simplicial complex consisting of all chains in $P$) and the $h$-polynomial of $\Delta(P)$ is defined as $h(\Delta(P),t)=\sum_{F\in\Delta(P)}t^{\#F}(1-t)^{d+1-(\#F)}$, where $d$ is the rank of $P$.
Proof of Proposition 3.2.
Let us write $\tau={\rm rev}(\sigma)$. By Equation (12) applied to ${\mathcal{Q}}_{\sigma}$ and since
$\sum_{m\geq 0}{\mathcal{Z}}(P_{\tau},m+1)t^{m}=\frac{h(\Delta(P_{\tau}),t)}{(1-t)^{n+1}}$
(see, for instance, [2, Equation (8)]) and ${\rm Ehr}({\mathcal{Q}}_{\sigma},t)={\mathcal{Z}}(P_{\tau},t+1)$ by Theorem 1.3, we have $h^{\ast}({\mathcal{Q}}_{\sigma},t)=h(\Delta(P_{\tau}),t)$. Thus, it suffices to show that $h(\Delta(P_{\tau}),t)$ is equal to the sum on the right-hand side of Equation (13).
For that reason, we consider the following edge labeling of the poset $\widehat{P}_{\tau}=P_{\tau}\cup\{\hat{1}\}$, obtained from $P_{\tau}$ by adding a maximum element $\hat{1}$. We recall that $P_{\tau}$ is the subposet of the product order ${\mathbb{N}}^{n}$ consisting of all points $(a_{1},a_{2},\dots,a_{n})\in{\mathbb{N}}^{n}$ such that
$a_{s_{i}+1}+a_{s_{i}+2}+\cdots+a_{n}\leq r_{i+1}+r_{i+2}+\cdots+r_{k}=n-s_{i}$
for $0\leq i<n$. The cover relations are the pairs $(u,v)\in P_{\tau}\times P_{\tau}$ such that $v$ can be obtained from $u$ by adding 1 in one of its coordinates. Given such a pair we set $\lambda(u,v)=i$, if $u$ and $v$ differ in the $i$th coordinate. We also set $\lambda(u,\hat{1})=1$ for every maximal element $u\in P_{\tau}$ and claim that the resulting edge labeling $\lambda:{\mathcal{E}}(\widehat{P}_{\tau})\rightarrow{\mathbb{Z}}$ is an EL-labeling. Given a nonsingleton closed interval $[u,v]$ of $\widehat{P}_{\tau}$, we need to verify that there is a unique rising (with respect to $\lambda$) maximal chain in $[u,v]$ and that this chain is lexicographically least among all maximal chains of $[u,v]$. This holds in case $v\neq\hat{1}$ because then the labels of the maximal chains of $[u,v]$ are exactly the permutations of a multiset; it holds if $v=\hat{1}$ as well since then there is a unique maximal chain with label $(1,1,\dots,1)$ and this chain is the only rising maximal chain of $[u,v]$.
Since $\lambda$ is an EL-labeling, by Equation (14) we have
$h(\Delta(P_{\tau}),t)=\sum_{w}t^{{\rm des}(w)},$
where $w=(w_{n},\dots,w_{1},w_{0})$ ranges over the labels of all maximal chains of $\widehat{P}_{\tau}$ and ${\rm des}(w)$ is the number of indices $i\in[n]$ such that $w_{i}>w_{i-1}$. We observe that $w$ is such a label if and only if $w_{0}=1$ and at least $s_{i}$ entries of $(w_{n},\dots,w_{2},w_{1})$ are less than or equal to $s_{i}$ for every $i\in[k]$ and the proof follows. ∎
Remark 3.3.
Proposition 3.2 specializes to a well known combinatorial interpretation of the $h^{\ast}$-polynomial (equivalently, of the $h$-polynomial of the $n$-fold edgewise subdivision) of the standard $n$-dimensional simplex when $\sigma=(n)$ has one part; see, for instance, [8, Corollary 8] for a generalization and refinement. A generalization of Proposition 3.2, valid for the polytopes in ${\mathbb{R}}^{n}$ defined by the equations
$x_{1}+x_{2}+\cdots+x_{s_{i}}\leq c_{1}+c_{2}+\cdots+c_{i}$
for $1\leq i\leq k$, where $c_{1},c_{2},\dots,c_{k}$ are positive integers, has been found by Xue Yan [14].